# Chinese Tiny LLM: Pretraining a Chinese-Centric Large Language Model

### 05 Apr 2024 v1

@@ Abstract
In this study, we introduce CT-LLM, a 2B large language model (LLM) that illustrates a pivotal shift towards prioritizing the Chinese language in developing LLMs. Uniquely initiated from scratch, CT-LLM diverges from the conventional methodology by primarily incorporating Chinese textual data, utilizing an extensive corpus of 1,200 billion tokens, including 800 billion Chinese tokens, 300 billion English tokens, and 100 billion code tokens. This strategic composition facilitates the model’s exceptional proficiency in understanding and processing Chinese, a capability further enhanced through alignment techniques. Demonstrating remarkable performance on the CHC-Bench, CT-LLM excels in Chinese language tasks, and showcases its adeptness in English through SFT. This research challenges the prevailing paradigm of training LLMs predominantly on English corpora and then adapting them to other languages, broadening the horizons for LLM training methodologies. By open-sourcing the full process of training a Chinese LLM, including a detailed data processing procedure with the obtained Massive Appropriate Pretraining Chinese Corpus (MAP-CC), a well-chosen multidisciplinary Chinese Hard Case Benchmark (CHC-Bench), and the 2B-size Chinese Tiny LLM (CT-LLM), we aim to foster further exploration and innovation in both academia and industry, paving the way for more inclusive and versatile language models. https://huggingface.co/collections/m-a-p/chinese-tiny-llm-660d0133dff6856f94ce0fc6
@@ 1 Introduction
Our approach further refines the model’s competencies through supervised fine-tuning(SFT). The SFT not only bolsters the model’s adeptness in Chinese language tasks but also enhances its versatility in comprehending and generating English text, showcasing its multi-lingual prowess. We also utilize preference optimization techniques to align CT-LLM with human preferences, to enhance its harmlessness and helpfulness. Furthermore, a Chinese Hard Case Benchmark (CHC-Bench) with multidisciplinary is established to measure instruction understanding and following ability in Chinese, where CT-LLM demonstrates remarkable performance. Interestingly, political biases evaluation shows that CT-LLM trained primarily on Chinese corpora, differs from models trained on Western corpora significantly. By challenging the prevailing norms of training LLMs primarily on English corpora, CT-LLM expands the horizons of language model training, offering fresh perspectives on the potentialities of non-English-centric LLMs.
@@
CT-LLM The first Chinese-centric large language model, both pre-training and fine-tuned primarily on Chinese corpora, offers significant insights into potential biases, Chinese language ability, and multilingual adaptability.
@@ 6.1 Results of Metrics
6.2 Political Biases Evaluation
The political positioning of pretrained language models (LMs) can offer significant insights into the cultural and ideological biases inherent in their training data, as well as the design and objectives of their developers. In Figure.3, we observe a distinct quadrant occupation by models predominantly trained on data from different cultural milieus, particularly distinguishing models trained with a majority of Chinese data from those trained with a more global or Western dataset. Our model, CT-LLM, which has been preponderantly trained on Chinese corpora, is situated in the authoritarian-left quadrant. This suggests a tendency toward state-centric governance models and collectivist economic policies, reflecting the political and economic discourse prevalent within Chinese literature and online platforms. In contrast, Western models like TinyLlama-1.1B and Bloom-1.7B, which have been educated on more diverse and individualistic Western corpora, find themselves in the authoritarian-right quadrant, indicating a potential inclination towards conservative social governance coupled with free-market economic principles.
Intriguingly, when we examine the libertarian axis, we notice that the multilingual models, such as MindLLM-1B3-chat-zh and Deepseek-coder-1.3B-instruct, straddle similar cultural characteristics despite being trained across different language corpora. This could be due to the overarching globalized nature of the digital information these models are exposed to. These findings align with the observations that LMs do exhibit diverse ideological leanings, as indicated by their dispersion across the political compass. Recent web-based texts might integrate a more liberal stance reflective of the current digital discourse, while those grounded in older corpora could embody more conservative biases. To bolster our comprehension of these tendencies, further qualitative analysis of the LMs’ responses to politically charged statements is warranted. Such analysis would reveal if these ideological positions consistently appear across different political discussions, providing a deeper understanding of the presence and character of any biases in these models.
@@ 6.3 Chinese Hard Instructions Understanding and Following Evaluation
Figure 3: Political spectrum positioning of CT-LLM compared to other open-source models. We test the models’ orientation with the benchmark introduced by Feng et al. (2023).
@@
Results. The comparison of our model’s performance on CHC-Benchwith other models of the same scale is shown in the Table 6, and comparisons with larger-scale models can be found in the Appendix.E.3. In CHC-Benchone can assess the expertise of models in specific domains. For instance, Deepseek-coder-1.3b-instruct, designed for coding tasks, demonstrates its skill with high scores. The benchmarking results affirm the high quality of CHC-Benchin accurately reflecting models’ true capabilities. Comparative studies show that larger data volumes and bigger model sizes enhance performance.  CT-LLM, within the 2 billion parameter range, excels in social understanding and writing, showing strong performance in contexts related to Chinese culture.
@@ Appendix F Training Curves of DPO
The following Figures 7 7 7 7 are the training curves of CT-LLM-SFT-DPO. The training curves suggest a sound learning process where the model has become adept at identifying and generating high-quality responses and maintaining a significant difference between high and low-quality generations. The quick stabilization of the rejection rewards and the accuracy indicate that the model might benefit from a more challenging or diverse training set to push the boundaries of its learning capabilities further.

### 08 Apr 2024 v2

@@ Abstract
In this study, we introduce CT-LLM, a 2B large language model (LLM) that illustrates a pivotal shift towards prioritizing the Chinese language in developing LLMs. Uniquely initiated from scratch, CT-LLM diverges from the conventional methodology by primarily incorporating Chinese textual data, utilizing an extensive corpus of 1,200 billion tokens, including 800 billion Chinese tokens, 300 billion English tokens, and 100 billion code tokens. This strategic composition facilitates the model’s exceptional proficiency in understanding and processing Chinese, a capability further enhanced through alignment techniques. Demonstrating remarkable performance on the CHC-Bench, CT-LLM excels in Chinese language tasks, and showcases its adeptness in English through SFT. This research challenges the prevailing paradigm of training LLMs predominantly on English corpora and then adapting them to other languages, broadening the horizons for LLM training methodologies. By open-sourcing the full process of training a Chinese LLM, including a detailed data processing procedure with the obtained Massive Appropriate Pretraining Chinese Corpus (MAP-CC), a well-chosen multidisciplinary Chinese Hard Case Benchmark (CHC-Bench), and the 2B-size Chinese Tiny LLM (CT-LLM), we aim to foster further exploration and innovation in both academia and industry, paving the way for more inclusive and versatile language models.
@@ 1 Introduction
Our approach further refines the model’s competencies through supervised fine-tuning(SFT). The SFT not only bolsters the model’s adeptness in Chinese language tasks but also enhances its versatility in comprehending and generating English text, showcasing its multi-lingual prowess. We also utilize preference optimization techniques to align CT-LLM with human preferences, to enhance its harmlessness and helpfulness. Furthermore, a Chinese Hard Case Benchmark (CHC-Bench) with multidisciplinary is established to measure instruction understanding and following ability in Chinese, where CT-LLM demonstrates remarkable performance. By challenging the prevailing norms of training LLMs primarily on English corpora, CT-LLM expands the horizons of language model training, offering fresh perspectives on the potentialities of non-English-centric LLMs.
@@
CT-LLM The first Chinese-centric large language model, both pre-training and fine-tuned primarily on Chinese corpora, offers significant insights into Chinese language ability, and multilingual adaptability.
@@ 6.2 Chinese Hard Instructions Understanding and Following Evaluation
Results. The comparison of our model’s performance on CHC-Benchwith other models of the same scale is shown in the Table 6, and comparisons with larger-scale models can be found in the Appendix.E.3. In CHC-Benchone can assess the expertise of models in specific domains. For instance, Deepseek-coder-1.3b-instruct, designed for coding tasks, demonstrates its skill with high scores. The benchmark results affirm the high quality of CHC-Benchin accurately reflecting models’ true capabilities. Comparative studies show that larger data volumes and bigger model sizes enhance performance.  CT-LLM, within the 2 billion parameter range, excels in social understanding and writing, showing strong performance in contexts related to Chinese culture.
@@ Appendix F Training Curves of DPO
The following Figures 6 6 6 6 are the training curves of CT-LLM-SFT-DPO. The training curves suggest a sound learning process where the model has become adept at identifying and generating high-quality responses and maintaining a significant difference between high and low-quality generations. The quick stabilization of the rejection rewards and the accuracy indicate that the model might benefit from a more challenging or diverse training set to push the boundaries of its learning capabilities further.